OpenStudy (anonymous):

Simplify: sin50 x cos20 - sin40 x sin20

OpenStudy (inkyvoyd):

Hey, do you mean sin(50)cos(20)-sin(40)sin(20?

OpenStudy (anonymous):

Yes

OpenStudy (inkyvoyd):

Please use asterisks (*) for multiplication. We usually use "x" for a variable.

OpenStudy (inkyvoyd):

hint: cos(x)=sin(90-x)

OpenStudy (anonymous):

cos(90-50)*cos20 - cos(90-40).cos(90-20) = -cos20 (cos(90-50) - cos (90-40))

OpenStudy (inkyvoyd):

cos(90-50)*cos20-cos(90-40)*cos(90-20) =cos(40)cos(20)-cos(50)cos(70)

OpenStudy (inkyvoyd):

=cos(40)cos(20)-sin(40)sin(20)

OpenStudy (inkyvoyd):

cos(alpha+beta)=cos(alpha)cos(beta)-sin(alpha)sin(beta)

OpenStudy (anonymous):

Thanks!